OpenStudy (rissa):

Find an equation of the line with x-intercept 2 and y-intercept -3.

OpenStudy (braydonlevi99):

Which type of equation do you want to use?

OpenStudy (rissa):

y=mx+b i assume

OpenStudy (braydonlevi99):

Yes, so you know that the x-intercept is 2. What does the x-intercept usually represent?

OpenStudy (rissa):

when the line intercepts at x (2,0)

OpenStudy (braydonlevi99):

X-Intercept can also mean the slope. And the y-intercept is usually your starting point. Now, plug in what you know.

OpenStudy (mathstudent55):

The x-intercept is where the line crosses the x-axis. It has nothing to do with slope.

OpenStudy (braydonlevi99):

What do you think the answer would be, Rissa?

OpenStudy (mathstudent55):

If you are given 2 points on a line, you can find the equation of the line. The x-intercept and the y-intercept are two points on the given line. That is enough information to find the equation of the line.

OpenStudy (mathstudent55):

@rissa The x-intercept is where the line crosses the x-axis. If the x-intercept is 2, what are the coordinates, (x, y), of the x-intercept?

OpenStudy (rissa):

(2,0) (0,-3)

OpenStudy (rissa):

than i find the slope ?

OpenStudy (rissa):

and plug in?

OpenStudy (braydonlevi99):



OpenStudy (braydonlevi99):

What is the slope of that line?

OpenStudy (braydonlevi99):

I'm pretty sure that this wold be your answer: \[y = \frac{ 3 }{ 2 } - 3\]

OpenStudy (mathstudent55):

Correct. Find the slope between the two intercepts.

OpenStudy (mathstudent55):

\(m = \dfrac{y_2 - y_1}{x_2 - x_1} = \dfrac{-3 - 0}{0 - 2} = \dfrac{-3}{-2} = \dfrac{3}{2}\)

OpenStudy (mathstudent55):

Now that you have the slope, m = 3/2, and the y-intercept, b = -2, you can use the slope-intercept form of the equation of a line which you wrote at the beginning of this post. y = mx + b Just substitute the values you have for m and b.

OpenStudy (rissa):

y=3/2x-3